# Aritro Pathak

### v2

@@ Abstract
$\overline{D}(E)=\limsup\limits_{l\to\infty}\frac{\log|E\cap[-l,l]^{2}|}{\log(2l)},\ D(E)=\limsup\limits_{||C||\to\infty}\frac{\log|E\cap C|}{\log||C||}.$ (1)
Here, $C\subset\mathbb{R}^{2}$ is any arbitrary cube with sides parallel to the axes, and $||C||$ denotes the length of each side of $C$.
In [9], the counting dimension was used in $\mathbb{Z}\subset\mathbb{R}$ to study the growth of certain subsets of $Z$ with zero upper Banach density. A natural Marstrand type projection theorem is proved there, with the counting dimension in this discrete setting resembling the Hausdorff dimension in the statement of the classical Marstrand projection theorem; see Theorem 1.2 in [9]. Later this was extended by Glasscock [5] who used the more general notion of the mass and counting dimension, in $\mathbb{Z}^{d}\subset\mathbb{R}^{d}$, and proved analogous projection theorems with the mass dimension as well.
@@
This is a tube of width 1. The line perpendicular to this tube and passing through the origin, has slope $u$. Henceforth we call this line $t_{\perp}$. The displacement of the point of intersection of the right edge of $t_{u,v}$ with $t_{\perp}$, along $t_{\perp}$ is given by the coordinate $v$ This is shown in the figure.
Figure 1: The tube $t_{u,0}$, with $v=0$ is shown, and the perpendicular line $t_{\perp}$.
@@
Let $E\subseteq\mathbb{R}^{2}$ be a $1$ separated set of mass dimension $\overline{D}(E)$. Then for all $v\in\mathbb{R}$, for Lebesgue-a.e. $u\in\mathbb{R}_{+}$,
@@
Let $E\subseteq\mathbb{R}^{2}$ be a $1$ separated set of mass dimension $\overline{D}(E)$. Then in the Lebesgue sense, for almost every tube $t_{u,v}$ of width $1$, slope $u$, and displacement $v$ along the projecting line, we have that $\overline{D}(E\cap t_{u,v})\leq\text{max}(0,\overline{D}(E)-1)$.
@@
The construction in the part (1) of this theorem is actually trivial; an analogous version theorem Theorem 3 is also true for the mass dimension, cf. Example 1 in Section 4 of [13]. This just consists of the set of the form $\{(x,y\in\mathbb{Z}^{2}:-a/2<x\leq a/2),y\geq 0\}$, for some $a\geq 2$.
@@
For any $\epsilon>0,u_{0}\in\mathbb{R}$, there exists $E\subseteq\mathbb{Z}^{2}$ , such that for all $(u,v)\in(-\epsilon,\epsilon)\times(-\epsilon,\epsilon)$,
@@
Terminology: Throughout the rest of the paper, we use the term ‘cluster’ of points to mean a subset of the form $T_{a,b}=\{(x,y)\in\mathbb{Z}^{2}:a_{1}\leq x\leq a_{2},b_{1}\leq y\leq b_{2}\}$ for some integers $a_{1}\leq a_{2},b_{1}\leq b_{2}$. We would also talk about clusters of points within a cone $\mathscr{C}$, or a cluster of points within a tube $t_{(u,v)}$ in which case it will be understood to mean $T_{a,b}\cap\mathscr{C}$, or $T_{a,b}\cap t_{(u,v)}$ respectively. When talking about such clusters, we will not enumerate the set of values of $a_{1},a_{2},b_{1},b_{2}$ as above, and it should be understood from the context that we mean a cluster of this type. Also, in several places, we use the term ’level’ to indicate the group of clusters of the set $E$ within the cone $C$, which are at approximately the same height from the origin (such as shown in Figure 2).
@@
Retain the setting of the cone of the previous paragraph. In order to construct a set $E$ supported within this cone $\mathscr{C}$, with counting dimension $D(E)>1$ such that we have a family of tubes $\{t_{(u,0)}\}$ for $u$ belonging in some small interval around the origin, with each tube in the family having an exceptionally high counting dimension greater than $(D(E)-1)$, we have to ensure that the set of points of $E$ do not accumulate in two dimensional clusters of growing length. In the examples below, we show how to construct a set $E$ that grows linearly. We can’t locate within this set any growing two dimensional sub-sequence that is growing to infinity. This is done while ensuring that each of the ray of tubes has a growing sub-sequence of points that ensures that it still has counting dimension 1. That would prove Theorem 4.
@@
Figure 2: The cone showing the clusters at any specific level of the construction of the set $E$, growing ’diagonally’ from one edge of the cone to the other. Every level of $E$ consists of such a diagonal sequence of points, and only one level of $E$ is shown in the figure. The width of each cluster is the same $w$, while the height of each cluster at the level $k$, is given by the sequence $n_{k}\to\infty$ as $k\to\infty$, which for simplicity can be chosen to be $n_{k}=k\ \forall k\in\mathbb{Z}$, as in the example.
@@
But the set $E$ so constructed is itself also of counting dimension exactly 1, since at each particular $m$’th level, we have a set of clusters, each of width $w(1+o(1))$, each of length $n_{1}+(m-1)$, the $k$’th one lying just above and to the right of the $(k-1)$’th cluster. If we had a square of length $K_{m}\cdot(n_{1}+(m-1))$ intersect the set $E$ at this appropriate height, then for the purpose of the counting dimension this is equivalent to having a vertical straight line of length $\approx K_{m}\cdot(n_{1}+(m-1))$ within this square. As $m\to\infty$, the lengths of these boxes go to infinity, and we have a set of counting dimension exactly 1. ∎
@@
We notice the similarity in the construction of the sets considered in Theorem 3(2) and Theorem 4, in that the growth of the points is in a diagonal manner in both cases. However, the set $E$ constructed in the proof of Theorem 3(2) has mass dimension 00 while the set constructed in the proof of Theorem 4 has mass dimension $1$.
@@
The construction follows from the previous example, by considering the same set $E$, withing the cone $C$. The set of tubes is parametrized by $\{(u,v):-\beta\leq u\leq\beta,-\epsilon\leq v\leq\epsilon\}$. With minor modifications to the argument in the previous theorem, we conclude that the set $E$ is such that the intersection of every tube with parameter in the range $(u,v)\in(-\beta,\beta)\times(-\epsilon,\epsilon)$, with $E$ has counting dimension exactly 1. The minor change here is that these tubes parameterized by the $u\neq 0$ parameters intersect the clusters of $E$ ’obliquely’, while still having counting dimension 1. ∎
Now the proof of Theorem 6 is also immediate with the same set $E$. This is the strongest statement we can make for the counting dimension.
@@
Again consider the same set $E$, within the same cone. Consider all the tubes $t_{(u,v)}$ whose right edges are lines that intersect the set $E$ with slopes in the range $(\cot\theta/2,\infty)$, and thus the slope $m(t_{\perp})$ of the line $t_{\perp}$ is so that $0\leq m(t_{\perp})\leq\tan(\theta/2)$ (The same is true of tubes whose right edges have negative slopes in the range $(-\cot(\theta/2),-\infty)$). In this case, any one of these tubes intersects all the ‘levels’ paramterized by large enough $k$, of the set $E$. In fact, it is clear that except for at most the first level it intersects, such a tube intersects both the top and bottom boundary layers of all the other levels. Thus, these tubes have clusters of points of increasingly large length, and these lengths go to infinity as the levels go to infinity. Thus all of these tubes have counting dimension exactly 1.
@@
    While we show that a positive Lebesgue measure set of tubes of width 1 have exceptionally large dimension, the question of constructing sets such that a set of broken lines $\{(x,y):y=\lfloor ux+v\rfloor\}$ with positive Lebesgue measure in the parameter space $(\tilde{u},\tilde{v})$, all have exceptionally large counting dimension, is not resolved here. The broken lines are sets that have vertical cross section 1, but the actual width is smaller than unity, and so there is no guarantee that a large number of broken lines have exceptionally high dimension, let alone a set of broken lines parametrized by a positive Lebesgue measure set in the $(u,v)$ space. Such a question will likely require a much finer study of the intersection patterns of tubes of arbitrary small width with the integer grid $\mathbb{Z}^{2}\subset\mathbb{R}^{2}$.

### v3

@@ Abstract
[…]
Here, $C\subset\mathbb{R}^{2}$ is any arbitrary cube with sides parallel to the axes, and $C||$ denotes the length of each side of $C$.
In [9], the counting dimension was used in $\mathbb{Z}\subset\mathbb{R}$ to study the growth of certain subsets of $\mathbb{Z}$ with zero upper Banach density. A natural Marstrand type projection theorem is proved there, with the counting dimension in this discrete setting resembling the Hausdorff dimension in the statement of the classical Marstrand projection theorem; see Theorem 1.2 in [9]. Later this was extended by Glasscock [5] who used the more general notion of the mass and counting dimension, in $\mathbb{Z}^{d}\subset\mathbb{R}^{d}$, and proved analogous projection theorems with the mass dimension as well.
@@
This is a tube of width 1. The line perpendicular to this tube and passing through the origin, has slope $u$. Henceforth we call this line $t_{\perp}$. The displacement of the point of intersection of the right edge of $t_{u,v}$ with $t_{\perp}$, is given by the coordinate $v$ . This is shown in the figure.
Figure 1: The tube $t_{u,0}$, with $v=0$ is shown, and the perpendicular line $t_{\perp}$. A tube $t_{u,v}$ with $v\neq 0$, is a parallel translation of $t_{u,0}$, along the line $t_{\perp}$.
@@
Let $E\subseteq\mathbb{R}^{2}$ be a $1$ separated set. Then for all $v\in\mathbb{R}$, for Lebesgue-a.e. $u\in\mathbb{R}_{+}$,
@@
Let $E\subseteq\mathbb{R}^{2}$ be a $1$ separated set. Then in the Lebesgue sense, for almost every tube $t_{u,v}$ of width $1$, slope $u$, and displacement $v$ along the projecting line, we have that $\overline{D}(E\cap t_{u,v})\leq\text{max}(0,\overline{D}(E)-1)$.
@@
The construction in the part (1) of this theorem is actually trivial; an analogous version Theorem 3 is also true for the mass dimension, cf. Example 1 in Section 4 of [13]. This just consists of the set of the form $\{(x,y\in\mathbb{Z}^{2}:-\epsilon/2<x\leq\epsilon/2),y\geq 0\}$.
@@
For any $\epsilon>0$, there exists $E\subseteq\mathbb{Z}^{2}$ , such that for all $(u,v)\in(-\epsilon,\epsilon)\times(-\epsilon,\epsilon)$,
@@
Notation: Throughout the rest of the paper, we use the term ‘cluster’ of points to mean a subset of the form $T_{a,b}=\{(x,y)\in\mathbb{Z}^{2}:a_{1}\leq x\leq a_{2},b_{1}\leq y\leq b_{2}\}$ for some integers $a_{1}\leq a_{2},b_{1}\leq b_{2}$. We would also talk about clusters of points within a cone $\mathscr{C}$, or a cluster of points within a tube $t_{(u,v)}$ in which case it will be understood to mean $T_{a,b}\cap\mathscr{C}$, or $T_{a,b}\cap t_{(u,v)}$ respectively. When talking about such clusters, we will not enumerate the set of values of $a_{1},a_{2},b_{1},b_{2}$ as above, and it should be understood from the context that we mean a cluster of this type. Also, in several places, we use the term ’level’ to indicate the group of clusters of the set $E$ within the cone $C$, which are at approximately the same height from the origin (such as shown in Figure 2).
@@
Retain the setting of the cone of the previous paragraph. In order to construct a set $E$ supported within this cone $\mathscr{C}$, with counting dimension $D(E)>1$ such that we have a family of tubes $\{t_{(u,0)}\}$ for $u$ belonging in some small interval around the origin, with each tube in the family having an exceptionally high counting dimension greater than (D(E)1)(D(E)-1), we have to ensure that the set of points of $E$ do not accumulate in two dimensional clusters of growing length. In the examples below, we show how to construct a set $E$ that grows linearly. We can’t locate within this set any growing two dimensional sub-sequence that is growing to infinity. This is done while ensuring that each of the ray of tubes has a growing sub-sequence of points that ensures that it still has counting dimension 1. That would prove Theorem 4.
@@
Figure 2: The cone showing the clusters at any specific level of the construction of the set $E$, growing ’diagonally’ from one edge of the cone to the other. Every level of $E$ consists of such a diagonal sequence of points, and only one level of $E$ is shown in the figure. The width of each cluster is the same $w$, while the height of each cluster at the level $k$, is given by the sequence $n_{k}\to\infty$ as $k\to\infty$, which for simplicity can be chosen to be $n_{k}=k,\ \forall k\in\mathbb{Z}$.
@@
But the set $E$ so constructed is itself also of counting dimension exactly 1, since at each particular $m$’th level, we have a set of clusters, each of width $w(1+o(1))$, each of length $n_{1}+(m-1)$, the $k$’th one lying just above and to the right of the (k1)(k-1)’th cluster. If we had a square of length $K_{m}\cdot(n_{1}+(m-1))$ intersect the set $E$ at this appropriate height, then for the purpose of the counting dimension this is equivalent to having a vertical straight line of length $\approx K_{m}\cdot(n_{1}+(m-1))$ within this square. As $m\to\infty$, the lengths of these boxes go to infinity, and we have a set of counting dimension exactly 1. ∎
@@
We notice the similarity in the construction of the sets considered in Theorem 3(2) and Theorem 4, in that the growth of the points is in a diagonal manner in both cases. However, the set $E$ constructed in the proof of Theorem 3(2) has mass dimension 0 while the set constructed in the proof of Theorem 4 has mass dimension $1$.
@@
The construction follows from the previous example, by considering the same set $E$, withing the cone $C$. The set of tubes is parametrized by $\{(u,v):-\beta\leq u\leq\beta,-\epsilon\leq v\leq\epsilon\}$. With minor modifications to the argument in the previous theorem, we conclude that the set $E$ is such that the intersection of every tube with parameter in the range $(u,v)\in(-\beta,\beta)\times(-\epsilon,\epsilon)$, with $E$, has counting dimension exactly 1. The minor change here is that these tubes parameterized by the $u\neq 0$ parameters intersect the clusters of $E$ ’obliquely’, while still having counting dimension 1. ∎
Now the proof of Theorem 6 is also immediate with the same set $E$.
@@
Again consider the same set $E$, within the same cone. Consider all the tubes $t_{(u,v)}$ whose right edges are lines that intersect the set $E$ with slopes in the range $(\cot\theta/2,\infty)$, and thus the slope $m(t_{\perp})$ of the line $t_{\perp}$ is so that $0\leq m(t_{\perp})\leq\tan(\theta/2)$ (The same is true of tubes whose right edges have negative slopes in the range $(-\tan(\theta/2),-\infty)$). In this case, any one of these tubes intersects all the ‘levels’ of $E$ parametrized by $k$ large enough. It is easy to see that except for at most the first level it intersects, such a tube intersects both the top and bottom boundary layers of all the other levels, and thus these tubes have counting dimension exactly 1.
@@
[…]